Musa is 60years old and joy is 25years old how many years from now will the product of their ages be 2144 years
Found 2 solutions by Theo, Edwin McCravy:
Answer by Theo(13342)

let x equal the number of years.

(60 + x) * (25 + x) = 2244

simplify to get 1500 + 60x + 25x + x^2 = 2244
combin like terms and reorder in descending order of degree to get:
x^2 + 85x + 1500 = 2244
subtract 2244 from both sides to get:
x^2 + 85x - 744 = 0
factor this quadratic equation to get:
(x - 8) * (x + 93) = 0
solve for x to get:
x = 8 or x = -93
x can't be negative, so solution is probably x = 8.

(60 + 8) * (25 + 8) = 68 * 33 = 2244.

8 years is your solution.







Answer by Edwin McCravy(20060)

There was a typo in your original post.  It originally read 2244 instead
of 2144. I assumed you meant 2144.


   (60+x)(25+x) = 2144
1500+60x+25x+x² = 2144
    x²+85x+1500 = 2144
     x²+85x-644 = 0
    (x-7)(x+92) = 0
     x-7 = 0; x+92 = 0
       x = 7     x = -92

In 7 years, they will be 67 and 32 and 67×32 = 2144

Also, it's laughable, but the -92 answer means that 92 years ago,
Musa was -32 (32 years before she was born) and Joy was -67,
(67 years before she was born), and the product of their "ages"
92 years before they were born is also +2144.  lol J

Edwin